OpenStudy (anonymous):

Solve for X x^2-4x-5 --------- x^2-8x+15

OpenStudy (anonymous):

Does this equal 0?

OpenStudy (anonymous):

No it does not

OpenStudy (anonymous):

(x+1)(x-5) --------- (x-3)(x-5) =(x+1) ------ (x-3) it is not for solve...it is to simplify.......

OpenStudy (anonymous):

if x^2 - 4x - 5 = x^2 - 8x + 15 then x = 5

OpenStudy (anonymous):

If you're just simplifying, it simplifies to: (x+1) ----- (x-3) If it is equal to 0, then x = -1